OpenStudy (anonymous):

3. Mr. S.P. Iderman needs a new suit. The suit costs $8,000. He can pay a down payment of $300 and then make a payment of $500 each month. If the interest rate is 6%, how much will Mr. Iderman owe at the end of the third month (after he makes three payments)?

OpenStudy (anonymous):

are you 9th grade connections academy?

OpenStudy (anonymous):

since he paid the down payment of $300 there still 8000-300 to pay = $7700 first month: paid $500 + 6% interest = $530 2nd month: paid $500 + 6% interest = $530 3rd month: paid $500 + 6% interest = $530 so Mr sp inderman paid a total of 1590 in the first 3 months so suit still costs 7700 and he paid 1590 so he need to pay $7700 - $1590 = $6,110 still needs to be paid

OpenStudy (anonymous):

thank you so much !!!!!